# High Dimensional Ensemble Kalman Filter

### v1

@@ 3 Consistency for fixed dimension case
The model error covariance matrix $\mathbf{Q}_{t}$, the observation error covariance matrix $\mathbf{R}_{t}$ and the initial analysis covariance matrix $\boldsymbol{\Sigma}_{0}^{a}$ satisfy $0<\varepsilon_{0}\leq\lambda_{\min}\left(\mathbf{Q}_{t}\right)\leq\lambda_{\max}\left(\mathbf{Q}_{t}\right)\leq 1/\varepsilon_{0}$, $0<\varepsilon_{0}\leq\lambda_{\min}\left(\mathbf{R}_{t}\right)\leq\lambda_{\max}\left(\mathbf{R}_{t}\right)\leq 1/\varepsilon_{0},$ and $\lambda_{\max}\left(\boldsymbol{\Sigma}_{0}^{a}\right)\leq 1/\varepsilon_{0},$ for a positive $\varepsilon_{0}$, respectively, where $\lambda_{\max}$ and $\lambda_{\min}$ are the maximum and minimum eigenvalue operators, respectively.
@@
Assumption 1 (i) and (ii) are two versions of the condition on the model operator $\mathcal{M}_{t}$. If model $\mathcal{M}_{t}$ is linear, it automatically satisfies both (i) and (ii) of Assumption 1. If $\mathcal{M}_{t}$ is nonlinear and Lipschitz continuous, $\mathcal{M}_{t}(\cdot)$ has at most linear growth at infinite interval, that is $\|\mathcal{M}_{t}(\mathbf{x})\|\leq\ell_{0}(1+\|\mathbf{x}\|)$ for any $\mathbf{x}\in R^{p}$ for a positive constant $\ell_{0}$. However, strongly nonlinear systems may not satisfy the Lipschitz condition. In that case, we assume $\mathcal{M}_{t}$ satisfies condition (ii), which to some extent is similar to the tangent linear hypothesis (Thepaut and Courtier, 1991; Bouttier and Courtier, 1999) in the 4D-Var, where $\mathcal{M}_{t}$ is determined in the linear approximation $\mathcal{M}_{t}(\mathbf{x}_{t}+\mathbf{h})\approx\mathcal{M}_{t}(\mathbf{x}_{t})+\dot{M}_{t}\mathbf{h}$ in the vicinity of an arbitrary state $\mathbf{x}_{t}$ for a perturbation $\mathbf{h}$, and the operator $\dot{M}_{t}$ is called the differential or the first derivative or tangent linear function of $\mathcal{M}_{t}$ at point $\mathbf{x}_{t}$. Assumption 2 assumes the eigenvalues of the three basic covariance matrices $\mathbf{Q}_{t}$, $\mathbf{R}_{t}$ and $\boldsymbol{\Sigma}_{0}^{a}$ are bounded away from zero and infinity. As presented in Lemma 3.1, under Assumption 2, the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ satisfies $\lambda_{\min}(\boldsymbol{\Sigma}_{t}^{f})\geq\epsilon_{0}>0$, which is the condition for the thresholding estimator for the high dimensional case in next section. With Assumptions 1 and 2, Lemma 3.1 shows that the eigenvalues of $\boldsymbol{\Sigma}_{t}^{f}$ are bounded from both ends, which assures $\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})\leq C$ for some positive constant $C$ and is also the condition for the bandable or the circular bandable covariance matrix classes. Assumption 3 assumes $\|\mathbf{H}_{t}\|\leq C$ for some positive constant $C$, which is a mild condition on the observation operator $\mathbf{H}_{t}$.
Lemma 3.1.
Suppose Assumption 2 holds, then the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ satisfies $\lambda_{\min}(\boldsymbol{\Sigma}_{t}^{f})\geq\epsilon_{0}>0$ for a positive $\epsilon_{0}$. If we further assume that Assumption 1 holds, then $0<\epsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma}_{t}^{f})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{t}^{f})\leq 1/\epsilon_{0}$ for a positive $\epsilon_{0}$.
@@
Proposition 3.1 suggests that the upper bound of the estimation error of the Kalman gain matrix estimation $\hat{\mathbf{K}}_{t}$ to $\mathbf{K}_{t}$ mainly depends on the bound of $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|$. Proposition 3.1 holds under weak conditions and actually is valid for both fixed and high dimensions. Note that for the fixed dimension data, the sample covariance matrix is consistent with the underlying population covariance at the $\sqrt{n}$ rate. As $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ is a sample covariance estimator, this means that $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|=O_{P}({n}^{-1/2})$ and thus $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|=O_{P}({n}^{-1/2})$.
Remark 1.
The last inequality is derived from $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|<1$ as $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}$ is the variance of $\mathbf{x}_{t}|\mathbf{y}_{1:t}$ and $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}<\boldsymbol{\Sigma}_{t}^{f}$ as long as $\mathbf{x}_{t}$ is correlated with $\mathbf{y}_{1:t}$. Let us take a close look at $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$. For the extreme case of $\mathbf{R}_{t}=\mathbf{0}$, $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ would be 00 if $\mathbf{H}_{t}$ is invertible. Intuitively, $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ would be close to 00 if $\mathbf{R}_{t}$ is close to $\mathbf{0}$. Suppose $p=q=1$, then $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|=(1+\mathbf{H}_{t}^{2}\sigma_{\boldsymbol{\Sigma}_{t}^{f}}^{2}/\sigma_{\mathbf{R}_{t}}^{2})^{-1}<1$. This means that $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ decreases with the increase of $\mathbf{H}_{t}^{2}\sigma_{\boldsymbol{\Sigma}_{t}^{f}}^{2}$ or decrease of $\sigma_{\mathbf{R}_{t}}^{2}$. For $p,q>1$, if $\mathbf{H}_{t}=\mathbf{I}$, then $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|=\|\mathbf{I}-\boldsymbol{\Sigma}_{t}^{f}(\boldsymbol{\Sigma}_{t}^{f}+\mathbf{R}_{t})^{-1}\|=\varepsilon_{0}^{2}/(1+\varepsilon_{0}^{2})$ if we assume $\mathbf{R}_{t}=\varepsilon_{0}^{2}\boldsymbol{\Sigma}_{t}^{f}$. This means that $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ will decrease with the decrease of $\varepsilon_{0}^{2}$, the ratio of $\mathbf{R}_{t}$ to $\boldsymbol{\Sigma}_{t}^{f}$.
In practice, the true observation error covariance matrix $\mathbf{R}_{t}$ may be unknown. Let $\hat{\mathbf{R}}_{t}$ be an estimator of $\mathbf{R}_{t}$ based on either knowledge or historical data. Corollary 3.1 below provides an upper bound of the estimation error of the $\hat{\mathbf{K}}_{t}$ to $\mathbf{K}_{t}$ when both the forecast and observation error covariance matrices $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{R}_{t}$ are estimated by $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ and $\hat{\mathbf{R}}_{t}$, respectively.
Corollary 3.1.
Under the assumptions of Proposition 3.1 and assume $\|\boldsymbol{\Sigma}_{t}^{f}\|\leq C$ for some positive constant $C$ and that $\hat{\mathbf{R}}_{t}$ is invertible with probability approaching $1$, then
$\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ $\leq\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|{\left\|\mathbf{H}_{t}\right\|}\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|+{\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})}{\|\mathbf{K}_{t}\|}\|\mathbf{R}_{t}-\hat{\mathbf{R}}_{t}\|$
$\leq{\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})}{\left\|\mathbf{H}_{t}\right\|}\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|+\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})\lambda_{\min}^{-1}(\mathbf{R}_{t})\left\|\mathbf{H}_{t}\right\|\|\boldsymbol{\Sigma}_{t}^{f}\|\|\mathbf{R}_{t}-\hat{\mathbf{R}}_{t}\|,$
where the Kalman gain matrix estimator takes the form of $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{\hat{R}}_{t}\big{)}^{-1}$.
Compared with the upper bound in Proposition 3.1, there is an extra term related to the bound of the estimation error of $\mathbf{R}_{t}$. Most of the results established in this paper are for the case of known $\mathbf{R}_{t}$ as is commonly assumed in the literature of the EnKF. Results for estimated $\mathbf{R}_{t}$ can be attained in the fashion shown in Corollary 3.1.
@@
Recall that $\hat{\mathbf{x}}_{t}^{a}={n}^{-1}\sum_{j=1}^{n}\mathbf{x}_{t,j}^{a}$ and $\hat{\mathbf{x}}_{t}^{f}={n}^{-1}\sum_{j=1}^{n}\mathbf{x}_{t,j}^{f}$ are, respectively, the analysis and forecast states of the EnKF. Based on the result in Proposition 3.1, we can establish the convergence of the EnKF forecast and analysis states $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\mathbf{x}}_{t}^{a}$ to their oracle Kalman Filter counterparts $\mathbf{x}_{t}^{f}$ and $\mathbf{x}_{t}^{a}$ based on the true $\boldsymbol{\Sigma}_{t}^{f}$ and the true underlying evolution given in (2.6)-(2.7). We also provide an upper bound for the difference between the forecast and the analysis states of the EnKF and the true state $\mathbf{x}_{t}$.
We first study the accuracy of the oracle $\mathbf{x}_{t}^{a}$ and $\mathbf{x}_{t}^{f}$ as “estimates” of the underlying true state $\mathbf{x}_{t}$. With (2.2), (2.3), (2.5), and (2.6), we have that $\mathbb{E}\big{\{}(\mathbf{x}_{t}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\big{\}}=\mathbf{0}$$\operatorname{Var}\big{\{}(\mathbf{x}_{t}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\big{\}}=\boldsymbol{\Sigma}_{t}^{f}$, $\mathbb{E}\left\{(\mathbf{x}_{t}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\right\}=\mathbf{0}$$\operatorname{Var}\left\{(\mathbf{x}_{t}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\right\}=\boldsymbol{\Sigma}_{t}^{a}$. Thus, $\mathbb{E}\|(\mathbf{x}_{t}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}=\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})\,\text{and }\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=\operatorname{tr}\left(\boldsymbol{\Sigma}_{t}^{a}\right)=\operatorname{tr}\big{(}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}\leq\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})$, and the discrepancies between the EnKF forecast and analysis states in comparison to the true state are, respectively,
$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t})|\mathbf{y}_{1:t-1}\|^{2}$ $=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}+\mathbb{E}\|(\mathbf{x}_{t}^{f}-\mathbf{x}_{t})|\mathbf{y}_{1:t-1}\|^{2}\,\text{ and}$ (3.1)
$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}$ $=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}+\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}.$
As mentioned before, the analysis state $\mathbf{x}_{t}^{a}$ of the Kalman Filter achieves optimality under the linear Gaussian state-space model (2.1). In fact, as shown in the following Proposition 3.2, under the linear model operator $\mathcal{M}_{t}$, Gaussian initial state and fixed dimensional $p$ and $q$, the asymptotic distributions of the forecast and analysis ensembles in the EnKF are the same as those of the Kalman Filter. Both the analysis state $\hat{\mathbf{x}}_{t}^{a}$ and the analysis covariance $\hat{\boldsymbol{\Sigma}}_{t}^{a}$ of the EnKF converge to the counterparts of the Kalman Filter $\mathbf{x}_{t}^{a}$ and $\boldsymbol{\Sigma}_{t}^{a}$, respectively.
Proposition 3.2.
Suppose Assumptions 2 and 3 hold, the model operator $\mathcal{M}_{t}$ is linear, $\mathbf{x}_{0}\mid\mathbf{y}_{0}\sim{\mathrm{N}}_{p}\left(\mathbf{x}_{0}^{a},\boldsymbol{\Sigma}_{0}^{a}\right)$, from which the initial analysis ensemble $\{\mathbf{x}_{0,j}^{a}\}_{j=1}^{n}$ are independently generated, then, for fixed dimensions $p$ and $q$, the forecast state of the EnKF obtained by (2.9) converges conditionally in probability or in distribution such that $\hat{\mathbf{x}}_{t}^{f}|\mathbf{y}_{1:t-1}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\mathbf{x}_{t}^{f}$, $\hat{\boldsymbol{\Sigma}}_{t}^{f}|\mathbf{y}_{1:t-1}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\boldsymbol{\Sigma}_{t}^{f}$, $\mathbf{x}_{t,j}^{f}|\mathbf{y}_{1:t-1}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathrm{N}_{p}(\mathbf{x}_{t}^{f},\boldsymbol{\Sigma}_{t}^{f})$, and the analysis state of the EnKF in (2.12) satisfies $\hat{\mathbf{x}}_{t}^{a}|\mathbf{y}_{1:t}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\mathbf{x}_{t}^{a}$, $\hat{\boldsymbol{\Sigma}}_{t}^{a}|\mathbf{y}_{1:t}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\boldsymbol{\Sigma}_{t}^{a}$, and $\mathbf{x}_{t,j}^{a}|\mathbf{y}_{1:t}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathrm{N}_{p}(\mathbf{x}_{t}^{a},\boldsymbol{\Sigma}_{t}^{a})$ as $n\rightarrow\infty$, where $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ and $\hat{\boldsymbol{\Sigma}}_{t}^{a}$ are the sample covariance matrices of the forecast ensemble $\{\mathbf{x}_{t,j}^{f}\}_{i=1}^{n}$ and the analysis ensemble $\{\mathbf{x}_{t,j}^{a}\}_{i=1}^{n}$, respectively.
Proposition 3.2 implies that with linear model operator $\mathcal{M}_{t}$ and Gaussian initial state, the asymptotic conditional distributions of the analysis ensemble $\{\mathbf{x}_{t,j}^{a}\}_{j=1}^{n}$ given $\mathbf{y}_{1:t}$ and the forecast ensemble $\{\mathbf{x}_{t,j}^{f}\}_{j=1}^{n}$ given $\mathbf{y}_{1:t-1}$ in the EnKF coincide with the distributions of $\mathbf{x}_{t}|y_{1:t}$ and $\mathbf{x}_{t}|y_{1:t-1}$ in the Kalman Filter, respectively.
The following Theorem 3.1 provides the upper bounds of the mean square error (MSE) between the analysis and forecasting states of the EnKF in comparison to those of the Kalman Filters according to the true underlying evolution given in (2.6)-(2.7) for the one-step assimilation for a general case where the model operator $\mathcal{M}_{t}$ is nonlinear.
To simplify notation, we denote $\Delta_{\mathbf{K},k}=\mathbb{E}\|\hat{\mathbf{K}}_{k}-\mathbf{K}_{k}\|^{2}$ and
Theorem 3.1.
Suppose Assumptions 1(i) or 1(ii), 2, 3 hold, and the forecast ensemble $\{\mathbf{x}_{t,j}^{f}\}_{j=1}^{n}$ are independently sampled from the distribution of $\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}$ with $\mathbf{x}_{t,j}^{f}\mid\mathbf{y}_{1:t-1}\sim\mathrm{N}(\mathbf{x}_{t}^{f},\boldsymbol{\Sigma}_{t}^{f})$ at time $t$. Then, the forecast and analysis states of the EnKF at time $t$ satisfy
$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}$ $={n}^{-1}\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})=O\big{(}{n}^{-1}\big{)},$
$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $=\operatorname{tr}\big{(}n^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}+\Omega_{t}=\operatorname{tr}\big{(}n^{-1}\boldsymbol{\Sigma}_{t}^{a}\big{)}+\Omega_{t}$ (3.3)
$=O\big{(}{n}^{-1}+\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}\big{)}=O\big{(}{n}^{-1}\big{)},$
where the results reported in the $O$-terms are valid for the fixed dimension case while the other results are valid for both fixed and high dimensional cases.
Note that the first term in (3.3) is the variance of the sample mean of the analysis ensemble using the true forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ and the second term $\Omega_{t}$ results from the estimation error of $\hat{\boldsymbol{\Sigma}}_{t}^{f}$. Theorem 3.1 implies that $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}=O\left(n^{-1}\right)=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ for the one-step data assimilation under the fixed dimensional case, which means both the forecast and analysis states of the EnKF converge to the underlying forecast and analysis states of the Kalman Filter with a convergence rate of $\sqrt{n}$ as $n\rightarrow\infty$. These are because the sample covariance matrix $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ is consistent with the underlying population covariance at the $\sqrt{n}$ rate, and it follows that $\Delta_{\mathbf{K},t}=\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}=O(n^{-1})$. The proof of these results can be found in the proof of Theorem 3.1 in the SM. It should also be noted that the above results are valid for the model operators $\mathcal{M}_{t}$ given in Assumption 1-(i) and (ii).
The discrepancies between $\hat{\mathbf{x}}_{t}^{f}$ or $\hat{\mathbf{x}}_{t}^{a}$ and the true state $\mathbf{x}_{t}$ are given by (3.1). As $\mathbf{x}_{t}$ is random with $(\mathbf{x}_{t}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\sim\mathrm{N}_{p}(\mathbf{0},\boldsymbol{\Sigma}_{t}^{f})$ in model (2.8), $\mathbb{E}\|(\mathbf{x}_{t}^{f}-\mathbf{x}_{t})|\mathbf{y}_{1:t-1}\|^{2}=\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})$ and $\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{a})$ are constants not related to the ensemble size $n$. This gives the result $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=O({n}^{-1}+1)=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t})|\mathbf{y}_{1:t-1}\|^{2}$, which is not related to the ensemble size $n$ in the leading order and can not converge to zero. This is due to the fact that the underlying state is a random variable. In contrast, $\mathbf{x}_{t}^{a}$ is a conditional mean and is possible to be estimated consistently.
The above justification on the validity of the Kalman Filter and EnKF algorithms is made for the fixed or low dimensional state vectors $\mathbf{x}_{t}$ where $p$ is of a smaller order of the ensemble size $n$ such that $p/n\to 0$. As the modern measurement technology on the earth system improves, the spatial resolution of the state variables becomes much higher. This makes $\mathbf{x}_{t}$ ultra high dimensional and creates high dimensional challenges on the validity of the EnKF, which is treated in the next section.
4 Convergence for High Dimensional Case
Remote sensing technology has been advancing the measurement of massive amounts of data sets for many geophysical processes, which produces high resolution data and brings high dimensional challenges. When the dimension $p$ of the state vector $\mathbf{x}_{t}$ is high so that $p\gg n$, the forecast ensemble sample covariance $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ in (2.10) is no longer consistent to the underlying covariance $\boldsymbol{\Sigma}_{t}^{f}$ as shown in Bai and Yin (1993). The inconsistency of the forecast error covariance estimator (2.10) to $\boldsymbol{\Sigma}_{t}^{f}$ is a major threat to the validity of the EnKF method. Although there are methods such as the localization and inflation approaches designed to rescue the EnKF in the high dimensional case, the statistical properties of these approaches remain largely unknown. The last 15 years have seen rapid development of high dimensional statistical inference, including efforts to construct consistent estimators for high dimensional covariance matrices which have proposed the banding, the thresholding and the tapering estimators (Bickel and Levina, 2008a, b; Cai, Zhang and Zhou, 2010).
In this section, we first introduce the covariance classes where consistent estimation of high dimensional covariance matrices can be made, and present several consistent estimators of the high dimensional forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$. Then, we provide results on the consistent estimation of the Kalman gain matrix $\mathbf{K}_{t}$ and evaluate the accuracy of the high dimensional assimilated state $\hat{\mathbf{x}}_{t}^{a}$ from the EnKF to the underlying $\mathbf{x}_{t}^{a}$ of the Kalman Filter.
4.1 Consistent estimation of high dimensional forecast error covariance matrix

### v2

@@ 3 Convergence of Ensemble Kalman Filter
The model error covariance $\mathbf{Q}_{t}$, the observation error covariance $\mathbf{R}_{t}$ and the initial analysis covariance $\boldsymbol{\Sigma}_{0}^{a}$ satisfy $0<\varepsilon_{0}\leq\lambda_{\min}\left(\mathbf{Q}_{t}\right)\leq\lambda_{\max}\left(\mathbf{Q}_{t}\right)\leq 1/\varepsilon_{0}$, $0<\varepsilon_{0}\leq\lambda_{\min}\left(\mathbf{R}_{t}\right)\leq\lambda_{\max}\left(\mathbf{R}_{t}\right)\leq 1/\varepsilon_{0},$ and $\lambda_{\max}\left(\boldsymbol{\Sigma}_{0}^{a}\right)\leq 1/\varepsilon_{0},$ for a positive $\varepsilon_{0}$, respectively, where $\lambda_{\max}$ and $\lambda_{\min}$ are the maximum and minimum eigenvalue operators, respectively.
@@
Assumption 1 (i) and (ii) are two versions of the condition on the model operator $\mathcal{M}_{t}$. If model $\mathcal{M}_{t}$ is linear, it automatically satisfies both (i) and (ii) of Assumption 1. If $\mathcal{M}_{t}$ is nonlinear and Lipschitz continuous, $\mathcal{M}_{t}(\cdot)$ has at most linear growth at infinite interval. However, strongly nonlinear systems may not satisfy the Lipschitz condition. In that case, we assume $\mathcal{M}_{t}$ satisfies condition (ii), which is related to the tangent linear hypothesis (Thepaut and Courtier, 1991; Bouttier and Courtier, 1999) in the 4D-Var, where $\mathcal{M}_{t}$ is determined in the linear approximation $\mathcal{M}_{t}(\mathbf{x}_{t}+\mathbf{h})\approx\mathcal{M}_{t}(\mathbf{x}_{t})+\dot{M}_{t}\mathbf{h}$ in the vicinity of $\mathbf{x}_{t}$ for a perturbation $\mathbf{h}$, and the operator $\dot{M}_{t}$ is called the differential or tangent linear function of $\mathcal{M}_{t}$ at point $\mathbf{x}_{t}$. Assumption 2 assumes the eigenvalues of the three basic covariances $\mathbf{Q}_{t}$, $\mathbf{R}_{t}$ and $\boldsymbol{\Sigma}_{0}^{a}$ are bounded away from zero and infinity. From Lemma S.3.1 of the SM, the eigenvalues of $\boldsymbol{\Sigma}_{t}^{f}$ satisfy the conditions for the thresholding and bandable(or circular bandable) covariance classes under Assumption 2 and Assumptions 1-2, respectively. Assumption 3 puts a mild condition on the observation operator $\mathbf{H}_{t}$.
@@
The last inequality is derived from $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|<1$ as $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}$ is the variance of $\mathbf{x}_{t}|\mathbf{y}_{1:t}$ and $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}<\boldsymbol{\Sigma}_{t}^{f}$ as long as $\mathbf{x}_{t}$ is correlated with $\mathbf{y}_{1:t}$. Proposition 3.1 suggests that the upper bound of the estimation error in $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ mainly depends on the bound of $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|$. Proposition 3.1 holds under weak conditions and actually is valid for both fixed and high dimensions. Note that under fixed dimension, the sample covariance is consistent with the underlying population covariance at the $\sqrt{n}$ rate, that is $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|=O_{P}({n}^{-1/2})$ and thus $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|=O_{P}({n}^{-1/2})$. The results for the high dimensions are shown later. In practice, the true observation error covariance matrix $\mathbf{R}_{t}$ may be unknown. Corollary S.3.1 of the SM provides an upper bound of the estimation error of the $\hat{\mathbf{K}}_{t}$ to $\mathbf{K}_{t}$ when both $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{R}_{t}$ are estimated by $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ and $\hat{\mathbf{R}}_{t}$, respectively.
@@
Based on the result in Proposition 3.1, we can establish the convergence of the EnKF forecast and analysis states $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\mathbf{x}}_{t}^{a}$ to their oracle Kalman Filter counterparts $\mathbf{x}_{t}^{f}$ and $\mathbf{x}_{t}^{a}$ based on the true $\boldsymbol{\Sigma}_{t}^{f}$ and the true underlying evolution given in (2.5)-(2.6) by firstly establishing the convergence of $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ to $\boldsymbol{\Sigma}_{t}^{f}$. For the fixed dimension case, the sample covariance matrix is consistent such that $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|=O_{P}({n}^{-1/2})$. However, when the dimension $p$ of the state vector $\mathbf{x}_{t}$ is high so that $p\gg n$, the forecast ensemble sample covariance $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ in (2.9) is no longer consistent to the underlying covariance $\boldsymbol{\Sigma}_{t}^{f}$ (Bai and Yin, 1993), which is a major threat to the validity of the EnKF method.
Next, we present several consistent estimators of the high dimensional forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ and their covariance classes before we give the convergence of the EnKF.
3.1 Consistent estimation of high dimensional $\boldsymbol{\Sigma}_{t}^{f}$
The banding estimator (Bickel and Levina, 2008b) of the forecast error covariance matrix with bandwidth $k$ is $B_{k}(\mathbf{S}_{t}^{f})=\big{[}s_{tij}^{f}\mathbf{1}(|i-j|\leq k)\big{]}$, where $0\leq k<p$ and $\mathbf{S}_{t}^{f}=[s_{tij}^{f}]_{p\times p}$ is the sample covariance of the EnKF forecast ensembles. The covariance matrix class suitable for the banding estimator is the following bandable covariance class, $\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\max_{j}\sum_{|i-j|>k}\left|\sigma_{ij}\right|\leq Ck^{-\alpha}\text{ for all }k>0\text{ and }0<\varepsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma})\leq\lambda_{\max}(\boldsymbol{\Sigma})\leq 1/\varepsilon_{0}\big{\}}$ for some positive $\alpha$, C and $\varepsilon_{0}$. This bandable pattern is generally consistent with the dependence between variables observed at sites $i$ and $j$ in geoscience studies, as the dependence becomes weaker as the distance between the two sites gets larger.
The banding estimator administrates a hard threshold as it sets all matrix elements beyond the $k$-th sub-diagonals to zero. Cai et al. (2010) updated the banding estimator with a tapering alteration which tapers down the level of the thresholding gradually
where $\mathbf{W}(k)=(w_{ij}(k))_{p\times p}$, with $w_{ij}(k)=(2/k)\left\{(k-|i-j|)_{+}-\left(k/2-|i-j|\right)_{+}\right\},$ is the tapering matrix at a tapering width $k$ which is usually an even integer and $1\leq k\leq p$.